OpenStudy (anonymous):

Which of the following equations is in standard form? A - 5x + 5y = 35 B - 2x – 9y = 15 C - –y = 2x D - y = –2x + 3

OpenStudy (shadowlegendx):

Standard Form: Ax + By = C

OpenStudy (shadowlegendx):

Which one looks like that?

OpenStudy (anonymous):

So it would be A right?

OpenStudy (shadowlegendx):

Correct

OpenStudy (anonymous):

Thanxx <3

jimthompson5910 (jim_thompson5910):

2x – 9y = 15 works as well since it is in the form Ax + By = C A = 2 B = -9 C = 15

jimthompson5910 (jim_thompson5910):

refresh the page if weird box symbols show up

OpenStudy (anonymous):

Thanxx Jim i put A and got it wrong. It ended up being B

jimthompson5910 (jim_thompson5910):

my guess is that they wanted A,B,C to be coprime (or relatively prime)

OpenStudy (anonymous):

Lol shadow played you

jimthompson5910 (jim_thompson5910):

so that means 5x + 5y = 35 5(x + y) = 35 x+y = 7 is the simplified form which is in Ax+By = C form

OpenStudy (anonymous):

Yes Shadow played me and I fell for it cx

OpenStudy (shadowlegendx):

Haha sorry I was debating between A and B, and it seems they wanted coprime. But why would they? @jim_thompson5910

OpenStudy (anonymous):

Ahahaa its okay cx

OpenStudy (shadowlegendx):

Both forms were correct, but behind the curtain they had some sort of preference

OpenStudy (anonymous):

I guess so but whatev it doesn't matter

OpenStudy (anonymous):

Riiiiiight @ShadowLegendX // check ur message thing @RosePetal_

jimthompson5910 (jim_thompson5910):

my guess is that they just want everything simplified/reduced fully

jimthompson5910 (jim_thompson5910):

and for A,B,C to remain integers

OpenStudy (triciaal):

it might have been an honest mistake this time for @ShadowLegendX to say correct at first I would think both A and B were correct

OpenStudy (shadowlegendx):

Oh, I see 5x + 5y = 35 is the same as 5(x + y) = 35 They just wanted it simplified. My teacher just says to show it in a form that represents it and he'll take that. So I guess it's just different learning styles

OpenStudy (anonymous):

Ohh okayy i get it now. They both look right now that i look back at it